An 80.0 kg climber climbs to the top of Mount Everest, which has a peak height of 8848 m. What is the climber’s potential energy with respect to sea level?
1
Expert's answer
2020-04-06T09:13:48-0400

The potential energy of climber with respect to sea level

"E_p=mgh=80.0\times 9.81\times 8848=6.94\times 10^6 \:\rm J=6.94\:\rm MJ."